OpenStudy (anonymous):

Solve the following for x. log2x+log3=log5

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif see rule number 1 there can you simplify the left-hand side?

OpenStudy (anonymous):

log 6x?

OpenStudy (anonymous):

\[\log2x = \log 5 - \log3\] \[\log2x = \log \frac{ 5 }{ 3 }\] now you can figure the rest out. I know you can do it!

OpenStudy (anonymous):

you could also approach it \[\log6x = \log5\]

OpenStudy (anonymous):

okay i got 0.737??

OpenStudy (jdoe0001):

yes is 6x, thus $$ \large { log(2x)+log(3) = log(5) \implies log(2x \times 3) = log(5)\\ \implies log(6x) = log(5)\\ \text{now usig the log cancellation rule of }\\ a^{log_ax} = x\\ log(6x) = log(5) \implies log_{10}(6x) = log_{10}(5)\\ 10^{log_{10}(6x)} = 10^{log_{10}(5)}\\ 6x = 5 } $$

OpenStudy (anonymous):

you should get 5/6

OpenStudy (anonymous):

ohh, i see what i did wrong! alright thank youuu

OpenStudy (jdoe0001):

yw